OpenStudy (anonymous):

Ohk @dan815 I need your help. I have an answer but i think its wrong. BUT ANYONE CAN STILL ADD TO THE HELP!!

OpenStudy (anonymous):

50 men 50 women. 20 men prefer baseball 30 prefer football. 25 women prefer watching baseball and 25 prefer football. If you pick someone who prefers baseball, what is the probability that they are a man?

OpenStudy (anonymous):

A. 1/5 B. 5/9 C. 1/4 D. 4/9

OpenStudy (anonymous):

now the answer i got was 1.1111111 but idk if thats right or not

OpenStudy (anonymous):

20+25=45 45/50=1.111111

OpenStudy (anonymous):

A 2X2 contingency table can be prepared as shown below. __________ MEN ___ WOMEN ___ TOTAL Prefer baseball ___ 20 _______ 25 _______ 45 Prefer football ____ 30 _______ 25 _______ 55 TOTAL _____50 ______ 50 _______ 100 It is a case of conditional probability. P( prefer football / woman) = P(women prefer football) / P(women)

OpenStudy (anonymous):

well i mean 50/45

OpenStudy (anonymous):

ohk let me try that really fast

OpenStudy (anonymous):

Sure. :)

OpenStudy (anonymous):

so its 25 right

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

= (25/100) / (50/100)

OpenStudy (anonymous):

Hold on. I dint something wrong.

OpenStudy (anonymous):

did*

OpenStudy (anonymous):

yea i was lost for a second on the 50/100

OpenStudy (anonymous):

so men that prefer baseball. 45 total right?

OpenStudy (anonymous):

25 men though

OpenStudy (anonymous):

50 men ina ll and only 20 prefer baseball

OpenStudy (anonymous):

in all*

OpenStudy (anonymous):

20 over 100 right because there is 100 people.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

20/100 = ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

simple division. not 5 that would be 5/1 or 100/20

OpenStudy (anonymous):

20/100 is what

OpenStudy (anonymous):

Just reduce it.

OpenStudy (anonymous):

but thats not an option 20/100 would be 0.2

OpenStudy (anonymous):

or what as a fraction?

OpenStudy (anonymous):

Reduce by 20

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

Divide 20 by 20 you get 1. divide 100/20 you get 5 = 1/5

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

thank you so much im sorry im slow sometimes

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

1/5 was wrong @DW101796